# Separating subsets from their images

### v1

@@ 7. Sharpness results for D
Let $G$ be a primitive permutation group of degree $n$. Then $\mathbf{m}(G)=\lceil(n+1)/2\rceil$ if, and only if, one of the following possibilities holds:
  • $n\geq 2$ and $G\cong\mathrm{Sym}(n)$ in natural action;
  • $n\geq 3$ and $G\cong\mathrm{Alt}(n)$ in natural action;
@@
  • $n=9$ and $G\cong{\rm AGL}_{1}(9)$, ${\rm PSU}_{3}(2)\cong M_{9}$, ${\rm A\Gamma L}_{1}(9)$, ${\rm ASL}_{2}(3)$, ${\rm AGL}_{2}(3)$, ${\rm PSL}_{2}(8)$, ${\rm P\Gamma L}_{2}(8)$;
@@
Let $G$ be an imprimitive permutation group of degree $n$ such that $\mathbf{m}(G)=\left\lceil(n+1)/2\right\rceil$. Then, $n$ is even, and $G$ defines a block system $\Sigma$ with either $n/2$ blocks of size $2$ or two blocks of size $n/2$. In particular, one of the following possibilities holds:
  1. $(a)$
    $n\leq 24$ and $G$ is permutationally isomorphic to a group TransitiveGroup(n,k)(n,k) from the library of transitive permutation groups in Magma [9] (based on [15, 23, 24]), where the pairs $(n,k)$ are recorded in Table 2;
  2. $(b)$
    $n\geq 16$ and the permutation group that $G$ induces on the block system $\Sigma$ is isomorphic to $\mathrm{Alt}(n/2)$ or $\mathrm{Sym}(n/2)$ in natural action (see Lemma 27 for the list of groups);
  3. $(c)$
    $n\geq 16$ and the permutation group that $G$ induces on a given block of $\Sigma$ isomorphic to $\mathrm{Alt}(n/2)$ or $\mathrm{Sym}(n/2)$ in natural action (see Lemma 28 and the previous discussion for the list of groups).
Degree $n$ # of groups Database numbers $k$
44 33 11, 22, 33
66 1919 22, 33, 55, 88, 99, $10$, $11$, $13$, $17$, $18$, 1919, 2121, $22$, $26$, $28$, $29$, $30$, $31$, $35$
88 $28$ 33, 99, $13$, $14$, $17$, $18$, 1919, 2121, $22$, $24$, $26$, $28$, $29$, $30$, $31$, $32$, $33$, $34$, $35$, $38$, $39$, $40$, $41$, $42$, $44$, $45$, $46$, $47$
$10$ $17$ $11$, $12$, $17$, 1919, 2020, $22$, $25$, $27$, $29$, $33$, $36$, $38$, $39$, $40$, $41$, $42$, $43$
$12$ $22$ $123$, $180$, $181$, $182$, $183$, $219$, $220$, $256$, $257$, $270$, $277$, $278$, $279$, $285$, $286$, $287$, $288$, $293$, $296$, $297$, $298$, $299$
$14$ $10$ $46$, $47$, $49$, $54$, $56$, $57$, $58$, $59$, $60$, $61$
$16$ $23$ $1078$, $1502$, $1506$, $1693$, $1798$, $1799$, $1801$, $1802$, $1803$, $1804$, $1805$, $1842$, $1843$, $1844$, $1860$, $1861$, $1878$, $1882$, $1883$, $1902$, $1903$, $1916$, $1940$
$18$ 66 $855$, $897$, $914$, $937$, $938$, $952$
$24$ 44 $18439$, $24652$, $24732$, $24924$
Table 2. List of imprimitive permutation groups of degree at most $24$ (with degree $14$ added) that meet the bound of D.
7.1. Primitive groups
We start our analysis with two general observations, that will greatly reduce the possibilities for primitive permutation group meeting the bound from D.
Lemma 23.
Let $G$ be a permutation group on $\Omega$ of degree $n$, let $c(g)$ be the number of cycles in the cyclic decomposition of $g\in G$, let $\ell(g)$ be the length of the shortest odd cycle in that composition (with $\ell(g)=0$ if $g$ contains no odd cycles), and for a non-negative integer $i$, let $\mathcal{O}_{i}$ be the set of elements $g$ in $G$ whose cycle decomposition contains at most $i$ odd cycles. Further, let $r(G)$ be the number of orbits of $G$ in its induced action on $\Omega\choose\lfloor n/2\rfloor$}. Suppose that $\mathbf{m}(G)=\lceil(n+1)/2\rceil$. Then, if $n$ is even,
@@ 7.1. Primitive groups
appearing in both formulas, for $n$ odd and for $n$ even, is just a restatement of the Cauchy-Frobenius Lemma (see [18, Theorem 1.7A]).
@@
First, we focus on the case when $n$ is even. Since we assume that every $(n/2)$-subset AA is self-separable, and there is a unique image of $A$ that is disjoint from it, it follows that $\left\{g\in G\mid A\cap A^{g}=\emptyset\right\}=G_{A}h$, where $h\in G$ is any element such that $A\cap A^{h}=\emptyset$. Hence,
@@
Given two functions $f,g:\mathbb{N}\to\mathbb{R}^{+}$, we write $f\sim g$ if the limit, as $n$ goes to infinity, of the quotient $f(n)/g(n)$ is $1$.
@@
$|G|\geq\frac{1}{2^{\lfloor n/2\rfloor}n}{n\choose\lfloor n/2\rfloor}\sim\sqrt{\frac{2}{\pi}}\cdot 2^{\frac{1}{2}(n-3\log_{2}n)}.$
@@
We are ready to characterize which primitive permutation groups meet the putative upper bound.
@@
By [32, Theorem 1.1], a primitive group whose degree exceeds $24$ is either a subgroup of $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$ with socle isomorphic to $\mathrm{Alt}(a)^{b}$, where the action of $\mathrm{Sym}(a)$ is on $k$-subsets (for some $1\leq k\leq n/2)$, the action of $\mathrm{Sym}(b)$ in on $b$ points, and the wreath product is endowed with the product action, or
@@
Using Wolfram Mathematica, we compute that the function appearing in Lemma 24 is greater that the bound from Equation G whenever $n$ is an even integer greater or equal to $116$, or $n$ is an odd integer greater or equal to $147$.
We have two cases to consider. If the degree of the permutation group $G$ exceeds $116$ or $147$, depending on the parity of its degree, by L, for every $a\geq 4$ and $b\geq 2$,
$\mathbf{m}(\mathrm{Sym}(a)\mathop{\mathrm{wr}}Sym(b))\leq\left\lceil\frac{a+1}{2}\right\rceil^{b}<\left\lceil\frac{a^{b}+1}{2}\right\rceil.$
In particular, if the degree exceeds $16$, we can assume that $G$ is of almost simple type. Moreover, Lemma 29 implies that $G$ is permutationally isomorphic to either $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$ endowed with their actions on $n$ points.
@@
Our search list get significantly shortened: the possible groups have degree at most $32$, with the only candidate of degree $32$ being ${\rm AGL}(5,2)$. We use a different approach for it. By randomly sampling subsets of size at most $16$ of the domain $\mathbb{F}_{2}^{5}$ and then testing them for existence of a group element that maps them to their complement, we found a number of sets of size $15$ that are not self-separable, thus showing that $\mathbf{m}({\rm AGL}(5,2))\leq 15$ (and we believe that $15$ is the exact value). The remaining groups all have degree $24$ or less, with the only groups of order $24$ being the Mathieu group $M_{24}$. These groups were dealt with an exhaustive search of $\lfloor(n+1)/2\rfloor$-subsets which are not self-separable. The groups where such a set was found were excluded form the list, and those that survived are precisely the ones listed in Lemma 21. Our code for these operations can be found in [5]. This concludes the proof. ∎
7.2. Imprimitive groups
We start this section with the following lemma, of independent interest, which gives a necessary condition for a permutation group to be $k$-homogeneus.
@@ 7.2. Imprimitive groups
Aiming for a contradiction, suppose that the claim of the lemma is false. Then the set
@@
is nonempty. Among all $(A,B)\in\mathcal{E}$ we choose one which minimises $|A\cap B|$. Observe that by our assumptions, the size $t$ of the intersection $A\cap B$ satisfies $1\leq t\leq k-1$. Thus,
@@
Observe that the sets $A\setminus B$ and $\Omega\setminus(A\cup B\cup C)$ are disjoint and of respective cardinalities ktk-t and n2kn-2k. Their union is therefore of size nktn-k-t. If $n-k-t\geq k$, then let $A^{\prime\prime}$ be an arbitrary $k$-subset of this union. Note that such a set $A^{\prime\prime}$ is disjoint from $A^{\prime}$, implying that there exist $g^{\prime}\in G$ such that $A^{\prime g^{\prime}}=A^{\prime\prime}$. Observe also that $B\cap A^{\prime\prime}=\emptyset$, implying that there exists $h\in G$ such that $A^{\prime\prime h}=B$. But then $A^{gg^{\prime}h}=B$, contradicting the fact that $(A,B)\in\mathcal{E}$. This contradiction implies that nkt<kn-k-t<k, or equivalently, 2k+tn=k(nkt)>02k+t-n=k-(n-k-t)>0. Since $2k\leq n-1$, it follows that $2k+t-n\leq t-1$, allowing us to choose a (2kn+t)(2k-n+t)-subset $D$ of $A\cap B$. Now let
@@
and observe that $|A^{\prime\prime}|=k$ and $A^{\prime\prime}\cap A^{\prime}=\emptyset$. Similarly as above, there exist $g^{\prime}\in G$ such that $A^{\prime g^{\prime}}=A^{\prime\prime}$. Observe also that $|B\cap A^{\prime\prime}|=|D|<t$, implying that there exists $h\in G$ such that $A^{\prime\prime h}=B$, which contradicts the fact that $(A,B)\in\mathcal{E}$. ∎

### v2

@@ 7.1. Primitive groups meeting the upper bound
Let $G$ be a primitive permutation group of degree $n$. Then $\mathbf{m}(G)=\lceil(n+1)/2\rceil$ if and only if one of the following holds:
  • $n\geq 2$ and $G\cong\mathrm{Sym}(n)$ in its natural action;
  • $n\geq 3$ and $G\cong\mathrm{Alt}(n)$ in its natural action;
@@
  • $n=9$ and $G\cong{\rm AGL}_{1}(9)$, ${\rm PSU}_{3}(2)$, ${\rm A\Gamma L}_{1}(9)$, ${\rm ASL}_{2}(3)$, ${\rm AGL}_{2}(3)$, ${\rm PSL}_{2}(8)$, ${\rm P\Gamma L}_{2}(8)$;
@@
The rest of the section is devoted to the proof of the above characterisation. We start our analysis with two general observations, that will greatly reduce the possibilities for primitive permutation group meeting the bound from E.
@@
Let $G$ be a permutation group acting on a set $\Omega$ of size $n$, let $c(g)$ be the number of cycles in the cyclic decomposition of $g\in G$, let $\ell(g)$ be the length of the shortest odd cycle in that composition (with $\ell(g)=0$ if $g$ contains no odd cycles), and for a non-negative integer $i$, let $\mathcal{O}_{i}$ be the set of elements $g$ in $G$ whose cycle decomposition contains at most $i$ odd cycles. Further, let $r(G)$ be the number of orbits of $G$ in its induced action on $\Omega\choose\lfloor n/2\rfloor$}. Suppose that $\mathbf{m}(G)=\lceil(n+1)/2\rceil$. Then, if $n$ is even,
@@
appearing in both formulas, for $n$ odd and for $n$ even, is just a restatement of the Cauchy-Frobenius Lemma (see [17, Theorem 1.7A]).
@@
First, we focus on the case when $n$ is even. By assumption, every $(n/2)$-subset AA is self-separable, and there is a unique image of $A$ that is disjoint from it. Hence, it follows that $\left\{g\in G\mid A\cap A^{g}=\emptyset\right\}=G_{A}h$, where $h\in G$ is any element such that $A\cap A^{h}=\emptyset$. Hence,
@@
Given two functions $f,g:\mathbb{N}\to\mathbb{R}^{+}$, we write $f\sim g$ if the limit of the quotient $f(n)/g(n)$, as $n$ goes to infinity, is $1$.
@@
$|G|\geq\frac{1}{2^{\frac{n-1}{2}}n}{n\choose\frac{n-1}{2}}\sim\sqrt{\frac{2}{\pi}}\cdot 2^{\frac{1}{2}(n-3\log_{2}n)}.$
@@
We are ready to characterise which primitive permutation groups meet the putative upper bound.
@@
By [29, Theorem 1.1], a primitive group whose degree exceeds $24$ is either a subgroup of $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$ with socle isomorphic to $\mathrm{Alt}(a)^{b}$, where the action of $\mathrm{Sym}(a)$ is on $k$-subsets (for some $1\leq k\leq a/2)$, the action of $\mathrm{Sym}(b)$ in on $b$ points, and the wreath product is endowed with the product action, or
@@
By elementary computations and by using the improved error term for Stirling formula proved in [37], we show that, for even integers $n\leq 300$,
$\frac{1}{2^{n/2}}{n\choose n/2}\geq n^{1+\log_{2}n},$
and that, for odd integers $n\leq 300$,
$\frac{1}{2^{\frac{n-1}{2}}n}{n\choose\frac{n-1}{2}}\geq n^{1+\log_{2}n}.$
Moreover, by checking a finite number of cases with a calculator, we can decrease these bounds even further. Indeed, the function appearing in Lemma 22 is greater that the bound from Equation G whenever $n$ is an even integer greater or equal to $116$, or $n$ is an odd integer greater or equal to $147$.
We have two cases to consider. If the degree of the permutation group $G$ exceeds $116$ or $147$, depending on the parity of its degree, by M, for every $a\geq 4$ and $b\geq 2$,
$\mathbf{m}(\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b))\leq\left\lceil\frac{a+1}{2}\right\rceil^{b}<\left\lceil\frac{a^{b}+1}{2}\right\rceil.$
In particular, if the degree exceeds $16$, we can assume that $G$ is of almost simple type. Moreover, Lemma 39 implies that $G$ is permutation isomorphic to either $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$ endowed with their actions on $n$ points.
@@
Our search list get significantly shortened: the possible groups have degree at most $32$, with the only candidate of degree $32$ being ${\rm AGL}(5,2)$. We use a different approach for it. By randomly sampling subsets of size at most $16$ of the domain $\mathbb{F}_{2}^{5}$ and then testing them for existence of a group element that maps them to their complement, we found a number of sets of size $15$ that are not self-separable, thus showing that $\mathbf{m}({\rm AGL}(5,2))\leq 15$ (and we believe that $15$ is the exact value). The remaining groups all have degree $24$ or less, with the only group of order $24$ being the Mathieu group $M_{24}$. These groups were dealt with an exhaustive search of $\lfloor(n+1)/2\rfloor$-subsets which are not self-separable. The groups where such a set was found were excluded form the list, and those that survived are precisely the ones listed in Lemma 20. Our code for these computations can be found in [5]. This concludes the proof. ∎
7.2. Imprimitive groups meeting the upper bound
Let us now move to imprimitive permutation groups $G$ meeting the upper bound on the parameter $\mathbf{m}(G)$ from E. A complete characterisation of these groups is given in Lemmas 31 and 37, which together with its primitive counterpart (Lemma 20) finishes the proof of E.
By J, the degree $n$ of an imprimitive permutation group $G$ with $\mathbf{m}(G)=\lceil(n+1)/2\rceil$ is even, and $G$ admits either a system of $n/2$ blocks of size $2$ or two blocks of size $n/2$. We shall consider these two cases in separately. We start our analysis with the following lemma of independent interest, which gives a necessary condition for a permutation group to be $k$-homogeneous.
@@ 7.2. Imprimitive groups meeting the upper bound
Aiming for a contradiction, suppose that the statement is false. Then the set
@@
is nonempty. Among all $(A,B)\in\mathcal{E}$, we choose one which minimises $|A\cap B|$. Observe that, by our assumptions, the size $t$ of the intersection $A\cap B$ satisfies $1\leq t\leq k-1$. Thus,
@@
Observe that the sets $A\setminus B$ and $\Omega\setminus(A\cup B\cup C)$ are disjoint and of cardinalities ktk-t and n2kn-2k, respectively. Their union is, therefore, of size nktn-k-t. If $n-k-t\geq k$, then let $A^{\prime\prime}$ be an arbitrary $k$-subset of this union. Note that such a set $A^{\prime\prime}$ is disjoint from $A^{\prime}$, implying that there exist $g^{\prime}\in G$ such that $A^{\prime g^{\prime}}=A^{\prime\prime}$. Observe also that $B\cap A^{\prime\prime}=\emptyset$, implying that there exists $h\in G$ such that $A^{\prime\prime h}=B$. But then $A^{gg^{\prime}h}=B$, contradicting the fact that $(A,B)\in\mathcal{E}$. This contradiction implies that nkt<kn-k-t<k, or equivalently, 2k+tn=k(nkt)>02k+t-n=k-(n-k-t)>0. Since $2k\leq n-1$, it follows that $2k+t-n\leq t-1$, allowing us to choose a (2kn+t)(2k-n+t)-subset $D$ of $A\cap B$. Now let
@@
and observe that $|A^{\prime\prime}|=k$ and $A^{\prime\prime}\cap A^{\prime}=\emptyset$. Similarly as above, there exist $g^{\prime}\in G$ such that $A^{\prime g^{\prime}}=A^{\prime\prime}$. Observe also that $|B\cap A^{\prime\prime}|=|D|<t$, implying that there exists $h\in G$ such that $A^{\prime\prime h}=B$, which contradicts the fact that $(A,B)\in\mathcal{E}$.
Hence $\mathcal{E}$ is empty, and the statement is proven. ∎